### v1

# Visualizing Linear Prediction

Abstract
Many statistics courses cover multiple linear regression, and present students with the formula of a prediction using the regressors, slopes, and an intercept. But is it really easy to see which terms have the largest effect, or to explain why the prediction of a specific case is unusually high or low? To assist with this the so-called grill plot is proposed. Its simplicity makes it easy to interpret, and it combines much information. Its main benefit is that it helps explainability of the linear formula as it is, without depending on how the formula was derived. The regressors can be numerical, categorical, or interaction terms, and the model can be linear or generalized linear. Another display is proposed to visualize correlations between predictors, in a way that is tailored for this setting.
Keywords: Effect size; Explainable prediction; Generalized linear model; Grill plot; Regression.
1 Introduction
Suppose someone has carried out a linear regression and tells you they predict the horsepower of a car by the formula
$-206.9+246.6*\mbox{topspeed}-13.1*\mbox{length}+0.063*\mbox{displacement}$
where topspeed is in miles per hour, length is in meters and engine displacement is in cubic centimeters. Which of these regressors has the biggest effect on the prediction? It is not enough to look for the largest coefficient. To resolve this, the typical approach is to first standardize each regressor by dividing it by its standard deviation. That works fine, but cannot be done for categorical regressors. We would like to assess the effect magnitude of regressors visually, in a way that includes categorical regressors. That is the topic of the next section.
A closely related issue is the question: Why is the prediction for a particular case so high or so low? An answer may be required when the prediction determines an important decision such as granting or denying a loan request, or treating a patient’s medical condition by surgery or medication. The ability to answer such a question is called explainability. It is often stressed that a regression tree is explainable because you can follow its branches, whereas most neural nets are not. Is a linear prediction explainable? On the one hand, a linear combination has a simple expression. On the other hand, how can you tell which of its terms were mainly responsible? This is a topic that deserves to be touched upon in a basic statistics class. We will show a visualization in Section 3, and look at correlations between predictors in Section 4.
In this note we are not concerned with the methodology used to obtain the fit. We will consider the prediction formula as a given, and study what it does. Our displays will not use the observed response variable either, or describe how well the prediction approximates it. That is addressed by other tools such as residual plots.
2 Spread and orientation of predictors
A linear prediction on numerical regressors $x_{1},\ldots,x_{p}$ is of the form
[…]
where $a$ is the intercept and the $b_{j}$ are slopes. We assume that the regressors have already been transformed if needed, and that outlying cases have been deleted or corrected. Let us now center the prediction terms, and denote them as $f_{j}:=b_{j}(x_{j}-\overline{x_{j}})$ The terms $f_{j}$ will be called predictors to distinguish them from the original regressors $x_{j}$ We can also compute the average prediction $\overline{\mbox{total\_pred}}$ which is sometimes called the centercept (Wainer, 2000), a name coined by John Tukey, because for the centered predictors it is the intercept. Next we also center the total prediction and denote it as $f:=\mbox{total\_pred}-\overline{\mbox{total\_pred}}$  so that (1) becomes
$f=\sum_{j=1}^{p}f_{j}$ (2)
without any intercept. In fact (2) is more general than (1) because a categorical variable also yields a predictor $f_{j}$ , that is obtained by coding its levels by binary dummies and adding up their predictions. We do not standardize the predictors $f_{1},\ldots,f_{p}$ so all terms are in the units of the total prediction. Finally, we measure the spread of each predictor $f_{j}$ by its standard deviation.
Our first example uses the Top Gear data from the R package robustHD (Alfons, 2016). It contains numerical and categorical variables about 297 cars, scraped from the website of the British television show Top Gear. We want to predict a car’s fuel efficiency from the type of fuel it uses (petrol or diesel), its weight in kilograms, the time in seconds it takes to accelerate from standstill to 60 miles per hour (accel), and the variable drive which has three levels: rear-wheel drive, front-wheel drive, and four-wheel drive (4WD). Following Henderson and Velleman (1981) who analyzed a similar dataset we predict gallons per mile (GPM), the inverse of miles per gallon (MPG).
Figure 1: A first attempt at visualizing the predictors $f_{1},f_{2},f_{3},f_{4}$ and the total prediction $f$ of (2).
After a standard regression analysis we have the predictors $f_{1}$ obtained from accel, $f_{2}$ from drive, $f_{3}$ from weight, $f_{4}$ from fuel, and the total prediction $f$ which by (2) is the sum of the four terms. Figure 1 plots all of them. We see that the $f_{j}$ and $f$ are centered, and they all have the same units of gallons per mile. The predictors $f_{1}$ and $f_{3}$ come from numerical regressors and take many values (here $n=297$), so their distributions are shown as histograms. The predictors $f_{2}$ and $f_{4}$ take very few distinct values, so they are represented by bar charts. We clearly see that the effect of predictor $f_{2}$ is quite small compared to that of $f_{1}$. One might have expected that the average of predictor $f_{4}$ which only takes two values would lie exactly in the middle of the two bars, but it doesn’t because the bars have different lengths, indicating that the positive prediction occurred more often than the negative one in this dataset.
Figure 1 already gives some intuition, but we can do more. It turns out that the weight predictor has the largest spread, followed by accel, fuel, and drive. In order to visualize this we can order the $f_{j}$ by decreasing spread. More importantly, we can label the ‘anonymous’ predictions in the scale of the original regressors, as in Figure 2. For the numerical regressors weight and accel we read off the values of $x_{j}$ without their slope coefficients. Predictions from the categorical variables fuel and drive are labeled by their levels. The vertical axis on the left measures the contribution of each term $f_{j}$ to the total prediction. For instance, increasing the weight of a car from the average to 2000 kg, while keeping everything else fixed, increases the prediction by about 0.005 gallons/mile. Replacing a diesel engine by a petrol engine raises the prediction by 0.008 gallons/mile. The effects of four-wheel drive and rear-wheel drive are about the same so their labels overlap, whereas front-wheel drive decreases the prediction.
Figure 2: Grill plot of the predictors in the Top Gear data.
Note that the weight line has an upward arrow and its histogram is shown in green, which indicates that increasing weight yields a higher prediction (again, if the other characteristics remain the same). On the other hand, longer acceleration times yield lower predictions, as reflected by the downward arrow and brown color. The lines of the categorical variables fuel and drive have no arrow because their levels are not ordered, and their distributions are shown in neutral grey. The same would happen for predictions from logical (Yes/No) variables and interaction terms. The rightmost line with the total prediction $f$ is labeled in its original scale that includes the centercept (which is about 0.026 here). Its range is a bit truncated in the plot, to put the focus on the predictors $f_{j}$ So we do not see its entire grey histogram, but that option is available.
We call Figure 2 a grill plot. With a little bit of imagination one can see the horizontal axis as a skewer which pierces some grilled pieces of food, both juicy numerical regressors and crispy categorical ones. The colors of the predictors evoke vegetables and meat, but can of course be changed by the user. An advantage of the grill plot over a list of predictor spreads is that we visualize the relative effects of the predictors. We would also notice it immediately if a predictor were highly skewed, and any outlier would stick out like a sore thumb.
The grill plot also works for generalized linear models, where the prediction (2) becomes
$f=g^{-1}\Big{(}\sum_{j=1}^{p}f_{j}\Big{)}.$ (3)
Here $g$ is a monotone link function. In logistic regression the link function is the logit function $g(p)=\log(p/(1-p))$, so (3) predicts a value between 0 and 1 using $g^{-1}(y)=1/(1+\exp(-y))$.
To illustrate this we analyze the well-known Titanic dataset which contains information on the 1309 passengers of the RMS Titanic. For its history and visualizations see Friendly et al. (2019). The data are freely available from https://www.kaggle.com/c/titanic/data and the R package classmap (Raymaekers and Rousseeuw, 2023). The binary response variable indicates whether the passenger survived or was a casualty. We want to predict survival by a logistic regression on the variables pclass, sex, age, sibsp, and parch. Here pclass is the cabin class, sex is M or F, and the passenger’s age is in years. The variables sibsp and parch count the number of siblings+spouses and parents+children aboard.
Figure 3: Grill plot of the logistic regression on the Titanic data.
Figure 3 shows the grill plot. It plots the $f_{j}$ and the total linear prediction $\sum_{j}f_{j}$ as before, but the labels on the rightmost axis are on the scale of the predicted survival probability $g^{-1}(\sum_{j}f_{j})$. This time we opted to plot estimated densities of age and the total prediction instead of their histograms. We see at first glance that the gender of the passenger is the most important variable, with predicted survival higher for females. This catches the eye to a similar extent as regression trees on this dataset making their first split on this variable, see e.g. Raymaekers and Rousseeuw (2022). Next up is the cabin class, followed by age which has a down arrow, meaning that younger people had a better chance at survival. Predictors sex and age reflect the motto ‘women and children first in the lifeboats,’ but being a first class passenger increases the prediction too. Clearly variable parch has only a tiny effect.
The grill plot is a member of the extensive family of parallel coordinate plots (Inselberg, 2009). It bears a slight resemblance to the so-called forest plot in meta-analysis, see e.g. Chang et al. (2022), which displays the outcomes of different studies with their confidence intervals. The grill plot contains no scientific novelty, but it is hoped that the way it visually combines much information provides helpful insight in the output of a regression. Figures 2 and 3 were obtained by the function grillplot in the Supplementary Material using the single R command grillplot(fit) where fit is an output object of the function lm() or glm(). The function grillplot figures out the rest. There are of course many options: the maximal number of predictors to be shown, whether to display density estimates instead of histograms, graphical parameters for colors and line widths, and so on. The formula of the regression may include logarithms of variables and interaction terms, such as $y\sim x_{1}+\log(x_{2})+x_{3}:x_{4}$ as illustrated in Section A of the Supplementary Material.
3 Why is the prediction for my case so high (low)?
The grill plot can also be used to explain the prediction of a single case, which may be one of the cases the model was fitted on (‘in-sample’), or a new case (‘out-of-sample’) for which we need a prediction. We illustrate this with another benchmark, the German credit dataset which is available from https://archive.ics.uci.edu/ml/datasets/statlog+(german+credit+data) and from the R package fairml (Scutari, 2023). It contains 1000 loan applications. We predict the success probability of a loan by logistic regression on the size of the loan (amount), its duration (months), the interest rate, the intended purpose of the loan (categorical with 10 levels), the number of clients responsible for paying back the loan (nclients), and the main client’s sex and age. Figure 4 shows the grill plot for the first case in the dataset.
Figure 4: Grill plot for case 1 of the German credit dataset.
In the grill plot we see that the purpose of the loan typically has the largest effect, followed by the three financial variables. The latter have downward arrows, so a longer duration, a higher rate, and a higher amount each decrease the predicted success probability. The variable age points upward, the model deems females more reliable than males, and being a one-client loan is considered positive. Now all the distributions are shown in grey, because the focus is on the predictor values. Those are red when the predictor value is above average, and blue if below average. Figure 4 is about a woman in her sixties requesting a small loan with short duration to buy a TV. Her predictor values are all above average except the one due to the high interest rate. Therefore the total prediction on the rightmost axis is very high, with estimated success probability over 90%.
Figure 4 suggests that the individual predictors are added, but does not actually show that. A variation of the plot that does visualize the addition is available as the option staircase = TRUE,illustrated in Figure 5 for case 2 of the dataset. In this plot the predictor values are added up by shifting each predictor up or down so its average aligns with the cumulative prediction to its left. Case 2 is a young man requesting a larger loan with longer duration, and his total prediction comes out rather low. If a loan is refused based on such a prediction, a graph like Figure 4 or Figure 5 facilitates explaining why to the client.
Figure 5: Grill plot for case 2 of the German credit dataset, in staircase style.
The choice between the basic style in Figure 4 and the staircase style in Figure 5 is a matter of taste. Whereas the latter shows the addition explicitly the plot also looks busier, and its appearance depends more strongly on the order of the individual predictors. The choice is left to the user. More examples of both styles are shown in the Supplementary Material.
While it can be very useful or even mandatory to be able to explain a prediction as in Figure 5, making the entire prediction equation (3) public may have the undesired side effect that people can game the system in favor of their desired outcome. After seeing Figure 5, the client might be tempted to increase his chances of getting a loan by indicating the intended purpose is to buy a used car because the u.car level is at the top, selecting a smaller number of months, and asking his mother to fill in the new application form to get more favorable values of age and sex, while keeping the other regressors unchanged. Running grillplot on this new set of regressors yields the figure in Section C of the Supplementary Material, in which the predicted probability becomes 89% instead of 48% .
4 Correlations between predictors
It is well-known that the coefficient of a regressor in multiple regression can have the opposite sign of its slope in a simple regression with that regressor alone. This often happens when that regressor is correlated with other regressors in the multiple model. Therefore it is useful to look at the correlations between regressors. One often visualizes these correlations by a heatmap, in which correlations of zero are shown in white, correlations close to 1 in dark red, and correlations close to -1 in dark blue, with a gradual change in between the extremes.
It is proposed to enrich such a display in two ways. First, by switching from regressors to predictors, so that also categorical regressors can be included. The left panel of Figure 6 shows such a display for the German credit data. For a numerical regressor, the sign of its slope coefficient matters. For a categorical regressor, its predictor is numerical so its correlation with other predictors exists.
Figure 6: Correlations between predictors of the German credit data. Left: in typical heatmap form. Right: here the side of a diagonal square is proportional to the spread of its predictor.
Note that this does not yet take the effect magnitude of the predictors into account. Our second modification is to plot the diagonal cells with sides that are proportional to the spread of their predictor. In this way the area of each diagonal cell is proportional to the variance of its predictor. The right panel of Figure 6 shows the result. We immediately see that the variable purpose has a big effect, whereas that of nclients is the smallest. Interestingly, such a correlation display with varying cell sizes did not appear to exist yet in the literature or on the internet, nor heatmaps of that type.
The off-diagonal part of the display consists of rectangles, which also have a meaningful interpretation. The area of such a rectangle is the product of the spreads of its predictors, and when we multiply that with the value of the correlation (represented by its color) we obtain the covariance between the predictors. So a dark color is more important in a larger rectangle (e.g. formed by months and amount) than in a smaller rectangle (e.g. formed by rate and amount). The proposed display thus depicts the covariance matrix between the predictors, which is relevant in a regression. Such a display is not recommended for a covariance matrix between variables in different units, but here all predictors are in the same units (those of the total prediction).
The R function corpred which produces this display has various options, such as sorting the predictors by decreasing spread, plotting the absolute values of the correlations, and making the area of the diagonal cells proportional to the spread of the predictors rather than their variance.
To illustrate what can happen when two regressors are highly correlated, we carry out a small experiment. We replace the variables months and nclients by their sum x1 = months + nclients and their difference x2 = months - nclients. Of course this is an artificial example, but it does yield two regressors with a correlation over $0.99$, and the same total prediction.
Figure 7: Grill plot of the German credit data when the regressors months and nclients are replaced by their sum x1 and their difference x2.
Figure 7 shows the resulting grill plot. It looks rather suspicious, because the artificial variables x1 and x2 have a wider spread than the total prediction on the right, which is the sum of these and other terms. Moreover x1 points downward and x2 points upward! They partly cancel each other. The correlation display in Figure 8 also looks unusual, with the big dark blue rectangles between x1 and x2. The correlation between x1 and x2 that was close to 1 has flipped sign between their predictors, again reflecting the canceling effect. In general, predictors are easiest to interpret when they only have small correlations with each other.
Figure 8: Correlations between predictors when the regressors months and nclients are replaced by their sum x1 and their difference x2.
5 Discussion
There are many other ways to measure the contribution of regressors. Verdinelli and Wasserman (2024) reviewed several, and found that there can be no ‘neutral’ way to do so, and that there is currently no ‘overall best’ way. The displays in this note may seem oversimplified but that makes them easy to interpret. Their main benefit is that they help explainability of the prediction rule as it is given, without depending on how it was derived. It would be possible to extend the displays to some related settings, but that falls outside the scope of this note.
Software availability. The R code and an example script that reproduces all the figures are available at https://wis.kuleuven.be/statdatascience/code/vlp_code_and_script.zip .
Acknowledgments. Mia Hubert and Jakob Raymaekers made helpful comments on the text.
Supplementary Material
Appendix A More figures of the Top Gear data
Figure 2 in the main text was obtained by the commands
fit = lm(1/MPG ~ accel + drive + weight + fuel + 0, data = cars)
main = "Top Gear Data"
grillplot(fit, main = main)
whereas the figures below were generated by
corpred(fit)
car = "Bentley Continental"
grillplot(fit, main = main, casetoshow = car, displaytype = "density",
          xlab = paste0("prediction for ", car))
car = "Kia Rio"
grillplot(fit, main = main, casetoshow = car, staircase = TRUE,
          xlab = paste0("prediction for ", car))
We see the correlations between the four predictors in the first display.
The Bentley Continental is a heavy car with high predicted GPM, whereas the Kia Rio is small and light and requires much less fuel:
To illustrate that the code can handle logarithms and interactions in the formula we run
[…]
This gets picked up automatically in the displays below. Interactions between categorical regressors, or between a numerical and a categorical variable, are visualized in the same way.
Appendix B More figures of the Titanic data
The correlation display below clearly indicates that the passenger’s sex has a large effect on the prediction, followed by passenger class and age. On the other hand, the effect of parch is minuscule. The off-diagonal colors are rather light, indicating smaller correlations between the Titanic predictors than between those of the Top Gear data.
On the next page we see the prediction for passenger 1, whose predicted survival probability is low. Most of the predictors are negative, mainly because the passenger was male and traveled in third class. The only positive predictor comes from his young age. Passenger 2 is a woman traveling in first class, so in spite of the slightly negative age-based predictor her total prediction is high.
Appendix C More figures of the German credit data
Here is the overall grill plot of the German credit data:
Next we look at a hypothetical new case, that is not in the original (training) dataset but was motivated at the end of Section 3. The new case is given as follows:
newc = list("u.car", 36, 2, 6000, 55, "F", 1)
names(newc) = c("purpose", "months", "rate", "amount", "age", "sex", "nclients")
gp = grillplot(fit, main = main, casetoshow = newc, staircase = TRUE)
This yields the plot on the next page:
In the output we see that the new predicted probability is about $89\%$ :
> print(gp\$casesummary, row.names = FALSE)
                          predictor    value
                             months -0.47190
                            purpose +1.02816
                             amount -0.25499
                               rate +0.23763
                                age +0.41640
                           nclients +0.03030
                                sex +0.14143
                                SUM +1.12701
                         centercept  0.95998
            TOTAL LINEAR PREDICTION  2.08699
 TOTAL PREDICTION IN RESPONSE UNITS  0.88963