### v2

# Explainable Linear and Generalized Linear Models by the Predictions Plot

Abstract
Many statistics courses cover multiple linear regression, and present students with the formula of a prediction using the input variables, slopes, and an intercept. But is it really easy to see which terms have the largest effect, or to explain why the prediction of a specific case is unusually high or low? To assist with this the so-called predictions plot is proposed. Its simplicity makes it easy to interpret, and it combines much information. Its main benefit is that it helps explainability of the prediction formula as it is, without depending on how the formula was derived. The input variables can be numerical or categorical; interaction terms are also handled, and the model can be linear or generalized linear. Another display is proposed to visualize correlations between prediction terms, in a way that is tailored for this setting.
Keywords: Effect size; Explainability; Linear prediction; Regression; Visualization.
1 Introduction
Suppose someone has carried out a linear regression and tells you they predict the horsepower (hp) of a car by the formula
$\widehat{hp}=2.466*\mbox{topspeed}-13.13*\mbox{length}+0.063*\mbox{displacement}-206.9$
where topspeed is in miles per hour, length is in meters and engine displacement is in cubic centimeters. (This example is worked out in the Supplementary Material.) Which of these input variables has the biggest effect on the prediction? It is not enough to look for the largest coefficient, because these depend on the units of the variables. To resolve this, the typical approach is to first standardize each input variable by dividing it by its standard deviation. That works fine, but cannot be done for categorical input variables. We would like to assess the effect magnitude of input variables visually, in a way that includes categorical input variables. That is the topic of the next section.
A closely related issue is the question: Why is the prediction for a particular case so high or so low? An answer may be required when the prediction determines an important decision such as granting or denying a loan request, or treating a patient’s medical condition by surgery or medication. The ability to answer such a question is called explainability. It is often stressed that a regression tree is explainable because you can follow its branches, whereas most neural nets are not. Is a linear prediction explainable? On the one hand, a linear combination has a simple expression. On the other hand, how can you tell which of its terms were mainly responsible? This is a topic that deserves to be touched upon in a basic statistics class. We propose a visualization in Section 3, and look at correlations between prediction terms in Section 4. We describe some related methods in Section 5, and Section 6 concludes.
In this note we are not concerned with the methodology used to obtain the fit. We will consider the prediction formula as a given, and study what it does. Our displays will not use the observed response variable either, or describe how well the prediction approximates it. That is addressed by other tools such as residual plots.
2 Spread and orientation of prediction terms
A linear prediction on numerical input variables $x_{1},\ldots,x_{p}$ is of the form
[…]
where $a$ is the intercept and the $b_{j}$ are slopes (coefficients). We assume that the input variables have already been transformed if needed, and that outlying cases have been deleted or corrected. Let us now center the terms with the $x_{j}$ , and denote them as $f_{j}:=b_{j}(x_{j}-\overline{x}_{j})$ The terms $f_{j}$ will be called prediction terms to distinguish them from the original input variables $x_{j}$ We can also compute the average total prediction $\overline{\mbox{total\_pred}}$ which is sometimes called the centercept (Wainer, 2000), a name coined by John Tukey, because for the centered input variables it is the intercept. Next we also center the total prediction and denote it as $f:=\mbox{total\_pred}-\overline{\mbox{total\_pred}}$  so that (1) becomes
$f=\sum_{j=1}^{p}f_{j}$ (2)
without any intercept. In fact (2) is more general than (1) because a categorical variable also yields a prediction term $f_{j}$ , that is obtained by coding its levels by binary dummies and adding up their predictions. We do not scale the prediction terms $f_{1},\ldots,f_{p}$ so all terms are in the units of the total prediction. Finally, we measure the spread of each prediction term $f_{j}$ by its standard deviation.
Our first example uses the Top Gear data from the R package robustHD (Alfons, 2016). It contains numerical and categorical variables about 297 cars, scraped from the website of the British television show Top Gear. We want to predict a car’s fuel efficiency from the time in seconds it takes to accelerate from standstill to 60 miles per hour (accel), the variable drive which has three levels: rear-wheel drive, front-wheel drive, and four-wheel drive (4WD), the car’s weight in kilograms, and the type of fuel it uses (petrol or diesel). Following Henderson and Velleman (1981) who analyzed a similar dataset we predict gallons per mile (GPM), the inverse of miles per gallon (MPG).
Figure 1: Top Gear data: an intermediate step toward visualizing the prediction term $f_{1}$ originating from the input variable accel, the term $f_{2}$ obtained from drive, $f_{3}$ from weight, and $f_{4}$ from fuel. On the right we see the total prediction $f$ of GPM computed as in (2).
After a standard regression analysis, briefly summarized in part A of the Supplementary Material, we can compute the prediction terms $f_{1}$ obtained from accel, $f_{2}$ from drive, $f_{3}$ from weight, $f_{4}$ from fuel, and the total prediction $f$ which by (2) is the sum of the four terms. Figure 1 is an intermediate step in the construction of the proposed display. It plots $f_{1}$, $f_{2}$, $f_{3}$, $f_{4}$ and $f$ next to each other. All the $f_{j}$ as well as $f$ are centered, and they all have the same units of gallons per mile. The prediction terms $f_{1}$ and $f_{3}$ come from numerical input variables and take many values, so their distributions are shown as histograms. The prediction terms $f_{2}$ and $f_{4}$ take very few distinct values, so they are represented by bar charts. We clearly see that the effect of prediction term $f_{2}$ is quite small compared to that of $f_{1}$. One might have expected that the average of prediction term $f_{4}$ which only takes two values would lie exactly in the middle of the two bars, but it doesn’t because the bars have different lengths, indicating that the higher prediction occurred more often than the lower one in this dataset.
Figure 1 already gives some intuition, but we can do more. It turns out that the weight prediction term has the largest standard deviation, followed by accel, fuel, and drive. In order to visualize this we can order the prediction terms $f_{j}$ by decreasing standard deviation. This is an example of effect ordering for data displays (Friendly and Kwan, 2003); see also (VanderPlas et al., 2023) for another instance of reordering variables. Moreover, we can label the ‘anonymous’ prediction terms in the scale of the original input variables, as in Figure 2. For the numerical input variables weight and accel we read off the values of $x_{j}$ without their slope coefficients. Predictions from the categorical variables fuel and drive are labeled by their levels. The vertical axis on the left measures the contribution of each term $f_{j}$ to the total prediction. For instance, increasing the weight of a car from the average (1562 kg) to 2000 kg, while keeping everything else fixed, increases the total prediction by about 0.005 gallons/mile. Replacing a diesel engine by a petrol engine raises the prediction by 0.008 gallons/mile. The effects of four-wheel drive and rear-wheel drive are about the same so their labels overlap, whereas front-wheel drive decreases the prediction.
Figure 2: Predictions plot of the Top Gear data. It shows the prediction terms originating from the input variables weight, accel, fuel and drive. The prediction terms are centered, and ordered by decreasing standard deviation. They are all in the units of the vertical axis on the left. The weight prediction term is shown in green with an upward arrow because its input variable has a positive coefficient, whereas accel is brown with a downward arrow due to its negative coefficient. The categorical variables have no arrow. On the right we see the total prediction in its units.
Note that the weight line has an upward arrow and its histogram is shown in green, which indicates that increasing weight yields a higher predicted fuel consumption (again, if the other characteristics remain the same). On the other hand, longer acceleration times yield lower predictions, as reflected by the downward arrow and brown color. The lines of the categorical variables fuel and drive have no arrow because their levels are not ordered, and their distributions are shown in neutral grey. The same will happen for predictions from logical (Yes/No) variables and interaction terms. The rightmost line with the total prediction $f$ is labeled in its original scale that includes the centercept (which is about 0.026 here). Its range is a bit truncated in the plot, to put the focus on the prediction terms $f_{j}$ So we do not see its entire grey histogram, but that option is available.
We call Figure 2 a predictions plot. The plural in “predictions” indicates that several predictions are plotted together. The colors of the prediction terms can of course be changed by the user. An advantage of the predictions plot over a list of standard deviations of input variables is that it visualizes the relative effects of the input variables. We would also notice immediately if a variable were highly skewed, and any outlier would stick out like a sore thumb.
The predictions plot also works for generalized linear models, where the prediction (2) becomes
$f=g^{-1}\Big{(}\sum_{j=1}^{p}f_{j}\Big{)}.$ (3)
Here $g$ is a monotone link function. In logistic regression the link function is the logit function $g(p)=\log(p/(1-p))$, so (3) predicts a value between 0 and 1 using $g^{-1}(y)=1/(1+\exp(-y))$.
To illustrate this we analyze the well-known Titanic dataset which contains information on the 1309 passengers of the RMS Titanic. For its history and visualizations see Friendly et al. (2019). The data are freely available from https://www.kaggle.com/c/titanic/data and the R package classmap (Raymaekers and Rousseeuw, 2023). The binary response variable indicates whether the passenger survived or was a casualty. We want to predict survival by a logistic regression on the variables pclass, sex, age, sibsp, and parch. Here pclass is the cabin class, sex is M or F, and the passenger’s age is in years. The variables sibsp and parch count the number of siblings+spouses and parents+children aboard.
Figure 3: Predictions plot of the logistic regression on the Titanic data. The prediction terms are again on the linear scale of the vertical axis on the left, but the labels of the total prediction on the right are not equally spaced since they indicate the predicted probability of survival. Here the continuous distributions are displayed by estimated densities instead of histograms.
Figure 3 shows the predictions plot. It plots the $f_{j}$ and the total linear prediction $\sum_{j}f_{j}$ as before, but the labels on the rightmost axis are on the scale of the predicted survival probability $g^{-1}(\sum_{j}f_{j})$. This time we opted to plot estimated densities of age and the total prediction instead of their histograms. We see at first glance that the gender of the passenger is the most important variable, with predicted survival higher for females. This catches the eye to a similar extent as regression trees on this dataset making their first split on this variable, see e.g. Raymaekers and Rousseeuw (2022). Next up is the cabin class, followed by age which has a down arrow, meaning that younger people had a better chance at survival. The prediction terms sex and age reflect the motto ‘women and children first in the lifeboats,’ but being a first class passenger increases the prediction too. Clearly variable parch has only a tiny effect.
It is hoped that the way the predictions plot visually combines much information provides helpful insight in the output of a regression. Figures 2 and 3 were obtained by the function predsplot in the Supplementary Material using the single R command predsplot(fit) where fit is an output object of the function lm() or glm(). The function predsplot figures out the rest. Of course there are many options: the maximal number of prediction terms to be shown, graphical parameters for colors and line widths, and so on. There is also an argument to specify whether to display histograms or estimated densities. For densities there are arguments that input the desired bandwidth as well. Here we used the defaults of stats::density(). The formula of the regression may include logarithms of variables and interaction terms, such as $y\sim x_{1}+\log(x_{2})+x_{3}:x_{4}$ as illustrated in Section A of the Supplementary Material.
3 Why is the prediction for my case so high (low)?
The predictions plot can also be used to explain the prediction of a single case, which may be one of the cases the model was fitted on (‘in-sample’), or a new case (‘out-of-sample’) for which we need a prediction. We illustrate this with another benchmark, the German credit dataset which is available from https://archive.ics.uci.edu/ml/datasets/statlog+(german+credit+data) and from the R package fairml (Scutari, 2023). It contains 1000 loan applications. We predict the success probability of a loan by logistic regression on the size of the loan (amount), its duration (months), the interest rate, the intended purpose of the loan (categorical with 10 levels), the number of clients responsible for paying back the loan (nclients), and the main client’s sex and age. Figure 4 shows the predictions plot for the first case in the dataset.
Figure 4: Predictions plot for case 1 of the German credit dataset. This loan application’s high predicted success probability on the right, over $0.9$ and shown in red, is explained by the mostly positive (red) values of its prediction terms, with only a single negative (blue) term. The distributions of the prediction terms are not colored, to put the focus on the colors and sizes of the prediction terms of the case being displayed.
In the predictions plot we see that the purpose of the loan has the largest effect, followed by the three financial variables. The latter have downward arrows, so a longer duration (‘months’), a higher rate, and a higher amount each decrease the predicted success probability. The variable age points upward, the model deems females more reliable than males, and being a one-client loan is considered positive. Now all the distributions are shown in grey, because the focus is on the prediction values. Those are red when the prediction value is above average, and blue if below average. From Figure 4 we can see that it is about a woman in her sixties requesting a small loan with short duration to buy a TV. Her prediction terms are all above average except the one due to the high interest rate. Therefore the total prediction on the rightmost axis is very high, with estimated success probability over 90%. It is also possible to add a profile curve to this plot, as illustrated in Section C of the Supplementary material.
Figure 4 suggests that the individual prediction terms are added, but does not actually show that. A variation of the plot that does visualize the addition is available as the option staircase = TRUE, illustrated in Figure 5 for case 2 of the dataset. In this plot the prediction terms are added by shifting each prediction term up or down so its average aligns with the cumulative prediction to its left. We see that case 2 is a young man requesting a larger loan with longer duration, and his prediction of credit worthiness comes out rather low. If a loan is refused based on such a prediction, a graph like Figure 4 or Figure 5 facilitates explaining why to the client.
Figure 5: Predictions plot for case 2 of the German credit dataset. In this ‘staircase style’ the prediction terms are added up from left to right, so after the last term we arrive at the total linear prediction. This particular loan application has many negative (blue) prediction terms, yielding a negative total linear prediction that translates to a relatively low success probability.
The choice between the basic style in Figure 4 and the staircase style in Figure 5 is a matter of taste. Whereas the latter shows the addition explicitly, the plot also looks busier, and its appearance depends more strongly on the order of the individual prediction terms. The choice is left to the user. More examples of both styles are shown in the Supplementary Material.
While it can be very useful or even mandatory to be able to explain a prediction as in Figure 5, making the entire prediction equation (3) public may have the undesired effect that people can game the system to obtain their favored outcome. After seeing Figure 5, the client might be tempted to increase his chances of getting a loan by indicating the intended purpose is to buy a used car because the u.car level is at the top, selecting a smaller number of months, and asking his mother to fill in the application form to get more favorable values of age and sex, while keeping the other input variables unchanged. Running predsplot on this new set of input variables yields the figure in Section C of the Supplementary Material, in which the predicted probability becomes 89% instead of 48% .
4 Correlations between prediction terms
It is well-known that the coefficient of an input variable in multiple regression can have the opposite sign of its slope in a simple regression with that input variable alone. This often happens when that input variable is correlated with other input variables in the multiple model. Therefore it is useful to look at the correlations between input variables. One often visualizes these correlations by a heatmap, in which correlations of zero are shown in white, correlations close to 1 in dark red, and correlations close to -1 in dark blue, with a gradual change in between the extremes.
We propose to enhance such a display in two ways. First, by switching from input variables to prediction terms, so that also categorical input variables can be included. The left panel of Figure 6 shows such a display for the German credit data. For a numerical input variable, the sign of its slope coefficient matters. For a categorical input variable, its prediction term is numerical so its correlation with other prediction terms exists.
Figure 6: Correlations between the prediction terms of the German credit data. The black squares on the diagonal have correlation 1. A red cell indicates a positive correlation, with color intensity proportional to the correlation value. A negative correlation is shown in blue, and zero correlation becomes white. The left panel is a customary correlation plot in which all cells have the same size. The right panel is the proposed version, where the side of a diagonal square is proportional to the standard deviation of its prediction term. The area of each off-diagonal cell is then proportional to the covariance of its prediction terms.
Note that this does not yet take the effect magnitude of the prediction terms into account. Our second modification is to plot the diagonal cells with sides that are proportional to the standard deviation of their prediction term. In this way the area of each diagonal cell is proportional to the variance of its prediction term. The right panel of Figure 6 shows the result. We immediately see that the variable purpose has a big effect, whereas that of nclients is the smallest. Interestingly, such a correlation display with varying cell sizes did not appear to exist yet in the literature or on the internet, nor heatmaps of that type.
The off-diagonal part of the display consists of rectangles, which also have a meaningful interpretation. The area of such a rectangle is the product of the standard deviations of its prediction terms, and when we multiply that with the value of the correlation (represented by its color) we obtain the covariance between the prediction terms. So a dark color is more important in a larger rectangle (e.g. formed by months and amount) than in a smaller rectangle (e.g. formed by rate and amount). The proposed display thus depicts the covariance matrix between the prediction terms, which is relevant in a regression. Such a display is not recommended for a covariance matrix between variables in different units, but here all prediction terms are in the same units (those of the total linear prediction).
The R function predscor which produces this display has various options, such as sorting the prediction terms by decreasing standard deviation, plotting the absolute values of the correlations, or making the area of the diagonal cells proportional to the standard deviation of the prediction terms rather than their variance.
To illustrate what can happen when two input variables are highly correlated, we carry out a small experiment. We replace the input variables months and nclients by their sum x1 = months + nclients and their difference x2 = months - nclients. Of course this is an artificial example, but it does yield two input variables with a correlation over $0.99$, and the same total prediction.
Figure 7: Illustration of the issue of highly correlated input variables, that is, near-multicollinearity. The input variables months and nclients of the German credit data are replaced by two artificial input variables x1 and x2 that have a large positive correlation between them. In this predictions plot their prediction terms point in opposite directions, indicating that their slope coefficients have opposite signs. These prediction terms are more spread out than the total linear prediction.
Figure 7 shows the resulting predictions plot. It looks rather suspicious, because the artificial variables x1 and x2 are more spread out than the total prediction on the right, which is the sum of these and other terms. Moreover, x1 points downward and x2 points upward! They partly cancel each other. The correlation display in Figure 8 also looks unusual, with the big dark blue rectangles between x1 and x2. The correlation between x1 and x2 that was close to 1 has flipped sign between their prediction terms, again reflecting the canceling effect. In general, prediction terms are easiest to interpret when they only have small correlations with each other.
Figure 8: Correlations between the prediction terms in Figure 7. The large diagonal squares labeled x1 and x2 reflect their high variance, and the dark blue cell between them indicates their large negative correlation.
5 Related methods
The predictions plot is a member of the extensive family of parallel coordinate plots (Inselberg, 2009). A related plot is the nomogram for binary logistic regression proposed by Zlotnik and Abraira (2015). It is also a parallel plot of prediction terms, but its purpose is computation rather than interpretation. The user needs to read off a score on each variable axis, add up the scores manually, and then look up the sum on another axis to obtain the resulting approximate probability. No distributions are visualized, and the predictions are not centered but instead their minima are aligned. This makes the nomogram hard to interpret, but interpretation was not its objective.
Effect plots (Fox, 2003) describe the effect of varying a single input variable on the total prediction. During this computation the other input variables are kept fixed at typical values, such as their average or median. A variation on this idea is to use the values taken by the remaining input variables in the actual dataset with $n$ cases, and then to average the $n$ results. Either way one obtains a plot of the total prediction as a function of the input variable being studied. The effect need not be linear, as the effect plot can be drawn for generalized linear models. The purpose of effect plots is similar to ours, to gain insight in the prediction, but the fact that it focuses on one input variable at a time makes it quite different from the approach presented here.
A coefficient plot, described for instance by Arel-Bundock (2022), does look at several input variables at the same time. But it does not plot the distributions of the input variables or the prediction. Instead it plots the regression coefficients associated with the input variables, that need to be numerical, as well as their confidence intervals. The coefficients are plotted as points, and the confidence intervals as line segments. The coefficients can be made comparable to each other if the input variables are first standardized, say to zero mean and unit standard deviation. The coefficient plot can summarize an entire model, but in a way very different from a predictions plot.
A forest plot, see for instance Chang et al. (2022), looks rather similar to a coefficient plot. But it displays the outcomes of different studies in a meta-analysis, with their confidence intervals. The predictions plot instead shows the distributions of the prediction terms in a single dataset.
6 Discussion
There are many other ways to measure the contribution of input variables. Verdinelli and Wasserman (2024) reviewed several, and found that there can be no ‘neutral’ way to do so, and that there is currently no ‘overall best’ way. The displays in this note may seem oversimplified but that makes them easy to interpret. Their main benefit is that they help explainability of the prediction rule as it is given, without depending on how it was derived. It would be possible to extend the displays to other settings, but that falls outside the scope of this note.
Software Availability. The R code with an example script that reproduces the figures is available at https://wis.kuleuven.be/statdatascience/code/predsplot_r_code.zip . It will be incorporated in an R package on CRAN.
Acknowledgments. Thanks go to Mia Hubert, Jakob Raymaekers, and the reviewers for helpful suggestions that improved the presentation.
Disclosure Statement. The author reports there are no competing interests to declare.
Supplementary Text
Appendix A More figures of the Top Gear data
We start with the example in the introduction, the fit
$\widehat{hp}=2.466*\mbox{topspeed}-13.13*\mbox{length}+0.063*\mbox{displacement}-206.9$
that was obtained from the standard regression
> fit = lm(hp ~ topspeed + length + displ, data = cars)
> summary(fit)
# Coefficients:
#               Estimate Std. Error t value Pr(>|t|)
# (Intercept) -2.069e+02  2.861e+01  -7.232 4.78e-12 ***
# topspeed     2.466e+00  1.373e-01  17.963  < 2e-16 ***
# length      -1.313e+01  6.187e+00  -2.122   0.0347 *
# displ        6.255e-02  2.808e-03  22.275  < 2e-16 ***
It is a simple example because all of the input variables are numeric. The question was, which of these has the biggest effect on the prediction. Using predsplot() we can easily answer this:
> gp = predsplot(fit, main = "Top Gear data")
> print(gp\$predsummary, row.names = F)
#         prediction term   stdev up/down
#                topspeed  68.380      up
#                  length   5.817    down
#                   displ  91.790      up
#  Total prediction of hp 149.200
We see that the prediction term from disp has the largest standard deviation, followed by topspeed, whereas length only has a tiny effect. The resulting predictions plot visualizes this:
> plot(gp\$p)
We clearly see that the range of the prediction term of length is tiny. This is confirmed by the new display of the correlations and standard deviations of the prediction terms:
> predscor(fit, sort.by.stdev = FALSE)
For Figure 2 in the main text we first ran a standard regression by
> fit = lm(GPM ~ accel + drive + weight + fuel, data = cars)
> summary(fit)
yielding, among other things:
Coefficients:
              Estimate Std. Error t value Pr(>|t|)
(Intercept)  1.681e-02  4.884e-03   3.443 0.000677 ***
accel       -1.356e-03  2.459e-04  -5.516 8.81e-08 ***
driveFront  -2.927e-03  1.619e-03  -1.808 0.071853 .
driveRear   -2.745e-04  1.577e-03  -0.174 0.861992
weight       1.131e-05  1.841e-06   6.143 3.25e-09 ***
fuelPetrol   8.427e-03  1.246e-03   6.761 9.98e-11 ***
The function lm() has encoded the categorical input variable drive with 3 levels by two binary dummy variables driveFront and driveRear, and the categorical variable fuel by the single binary dummy fuelPetrol. From this summary we cannot derive the relative effect of the four input variables accel, drive, weight and fuel.
For the numerical input variables accel and weight there is a known remedy, which is to divide them by their standard deviation and then rerun the fit:
> cars\$st.accel  = cars\$accel/sd(cars\$accel)
> cars\$st.weight = cars\$weight/sd(cars\$weight)
> st.fit = lm(GPM ~ st.accel + drive + st.weight + fuel, data = cars)
> summary(st.fit)
yielding
Coefficients:
              Estimate Std. Error t value Pr(>|t|)
(Intercept)  0.0168141  0.0048841   3.443 0.000677 ***
st.accel    -0.0043289  0.0007848  -5.516 8.81e-08 ***
driveFront  -0.0029267  0.0016189  -1.808 0.071853 .
driveRear   -0.0002745  0.0015774  -0.174 0.861992
st.weight    0.0044897  0.0007308   6.143 3.25e-09 ***
fuelPetrol   0.0084270  0.0012465   6.761 9.98e-11 ***
This changes the coefficients of acceleration and weight, all other coefficients remaining the same. From the absolute values of the new coefficients we see that the effect sizes of acceleration and weight are similar to each other, with that of weight being a bit higher. But we cannot do the same with a categorical variable like drive that is encoded by two dummies. Should we standardize both dummies, and even if we did, how can we combine their effects?
This is where the predictions plot comes in handy. We just run
> main = "Top Gear Data"
> gp = predsplot(fit, main = main)
> plot(gp\$p)
yielding Figure 2, and we get the standard deviations of the prediction terms by
> print(gp\$predsummary, row.names = F)
#          prediction term    stdev up/down
#                    accel 0.004329    down
#                    drive 0.001400
#                   weight 0.004490      up
#                     fuel 0.004104
#  Total prediction of GPM 0.009783
We see that the effect of the input variable drive is the smallest of the four.
The correlations between the four prediction terms are visualized by the line
> predscor(fit)
where we again see that the black square of drive is the smallest.
We can also make a predictions plot for a single case, in order to explain its prediction. To illustrate this we choose two rather extreme cars, the Bentley Continental and the Kia Rio:
> car = "Bentley Continental"
> plot(predsplot(fit, main = main, casetoshow = car, displaytype = "density",
                 xlab = paste0("prediction for ", car))\$p)
> car = "Kia Rio"
> plot(predsplot(fit, main = main, casetoshow = car, staircase = TRUE,
               xlab = paste0("prediction for ", car))\$p)
In the first plot we see that the Bentley gets a very high predicted gallons per mile (GPM). The prediction plot (this time with densities) explains its gas guzzling behavior by the fact that it is heavy (around 2300 kg as seen in the plot), accelerates from standstill to 60 miles per hour in under 5 seconds, runs on petrol, and has all-wheel drive (4WD).
The Kia Rio in the next plot finds itself in the opposite situation, as it is predicted to require very little fuel. The plot (this time in staircase style) explains this by the car’s low weight (around 1200 kg in the plot), slow acceleration, diesel engine, and front wheel drive.
To illustrate that the code can handle logarithms and interactions in the formula we run
[…]
This gets picked up automatically in the displays below. Interactions between categorical input variables, or between a numerical and a categorical variable, are visualized in the same way.
Appendix B More figures of the Titanic data
We now draw the correlation display of the Titanic logistic regression, by the lines
> fit <- glm(y ~ sex + age + sibsp + parch + pclass,family=binomial,data=titanic)
> predscor(fit, sort.by.stdev = FALSE)
It clearly indicates that the passenger’s sex has a large effect on the prediction, followed by passenger class and age. On the other hand, the effect of parch is minuscule. The off-diagonal colors are rather light, indicating smaller correlations between the Titanic prediction terms than between those of the Top Gear data.
The predictions plots for passengers 1 and 2 on the next page are made by
> plot(predsplot(fit, main = main, casetoshow=1)\$p)
> plot(predsplot(fit, main = main, casetoshow=2, staircase = TRUE)\$p)
In the first plot we see that the predicted survival probability of passenger 1 is low. Most of the prediction terms are negative, mainly because the passenger was male and traveled in third class. The only positive prediction term comes from his young age.
Passenger 2 is a woman traveling in first class, so in spite of the slightly negative age-based prediction term her total prediction is relatively high.
Appendix C More figures of the German credit data
Here is the overall predictions plot of the German credit data:
Next we look at a hypothetical new case, that is not in the original (training) dataset but was motivated at the end of Section 3. The new case is given as follows:
> newc = list("u.car", 36, 2, 6000, 55, "F", 1)
> names(newc) = c("purpose", "months", "rate", "amount", "age", "sex", "nclients")
> gp = predsplot(fit, main = main, casetoshow = newc, staircase = TRUE)
> plot(gp\$p)
This yields the plot on the next page:
In the output we see that the new predicted probability is about $89\%$ :
> print(gp\$casesummary, row.names = F)
                              prediction term    value
                                       months -0.47190
                                      purpose +1.02816
                                       amount -0.25499
                                         rate +0.23763
                                          age +0.41640
                                     nclients +0.03030
                                          sex +0.14143
                                          SUM +1.12701
                                   centercept  0.95998
            Total linear prediction of credit  2.08699
 Total prediction of credit in response units  0.88963
Note that non-staircase prediction plots for a single case, like Figure 4, can also be shown with a profile by the argument profile = TRUE :
> plot(predsplot(fit, main = main, casetoshow = 1,
                 displaytype = "density", profile = TRUE)\$p)
The profile is the faint grey broken line that connects the prediction terms of this case. It makes it easier to compare plots of this type made for different cases.